# Wall-Proximity Matters: Understanding the Effect of Device Placement with Respect to the Wall for Indoor Wi-Fi Sensing

### v2

@@ Abstract
Index Terms:
Wi-Fi sensing, channel state information (CSI), sensing coverage model, device placement, wall reflection.
@@ I Introduction
The extensive utilization of Wi-Fi for data communication in public areas has paved the way for a remarkable discovery in recent years—the untapped potential of Wi-Fi signals for sensing applications. Leveraging channel state information (CSI) to characterize the frequency response of wireless channels in Wi-Fi systems, researchers have unlocked new avenues for sensing capabilities [1, 2]. In indoor environments, Wi-Fi signals encounter reflections from a multitude of objects, including walls and human bodies, resulting in complex multi-path propagation. The variations in CSI values predominantly arise from the dynamic movements of humans. By leveraging time-frequency domain analysis, significant features can be extracted from CSI, enabling a diverse range of applications. Experimentally, CSI-based sensing has demonstrated remarkable capabilities in areas such as human activity recognition [3, 4], respiration monitoring [5, 6], crowd counting [7, 8], and indoor positioning [9, 10]. However, it is important to note that these applications have primarily been evaluated in controlled laboratory settings with carefully designed scenarios, and their real-life applicability and scalability are yet to be fully explored.
The real-world implementation of various applications in wireless systems has been limited due to several constraints. One major limitation is the instability of sensing performance, where CSI is highly susceptible to environmental factors such as interference and configurations [11, 12]. Consequently, changes in these factors have a direct impact on the accuracy performance, particularly in deep-learning-based sensing systems. Another limitation is the restricted sensing coverage of CSI-based applications. Despite Wi-Fi signals being capable of spanning significant distances, the accurate detection range is limited to just a few meters [13]. This limitation hampers the expansion of CSI-based applications into larger spaces and more complex scenarios. However, both limitations can be mitigated through a common solution: obtaining high-quality data. By employing reliable and precise data, the instability of sensing performance can be minimized, while also enabling the extension of sensing coverage.
How to obtain high-quality data in sensing systems? Besides employing dedicated signal processing schemes[13, 14, 15], we may still need to go back to the fundamental knowledge regarding the sensing coverage with respect to the locations of the transceivers. The placement of sensors and the design of the sensing network play a crucial role in the quality of collected data. The sensing coverage can be expanded by changing the locations of sensors [16]. Therefore, careful consideration of sensor placement and network design is vital for optimizing data collection, enhancing data quality, and addressing limitations such as unstable sensing performance and limited sensing coverage.
In prior research, the placement matters of wireless sensing have been investigated, leading to the observation that the sensing coverage is influenced by the distance between the transmitter (Tx) and receiver (Rx) [17]. However, this work primarily focuses on the impact of multipath on static paths and overlooks the influence of walls on dynamic paths in indoor environments. Another model addressing sensing coverage takes into account the beneficial effect of walls on through-wall sensing [18]. Nonetheless, this approach only considers scenarios where at least one device is positioned outside of the room, omitting indoor configurations. Therefore, while existing studies shed light on deployment issues, there remains a need to comprehensively address the dynamic effects of walls and consider diverse deployment scenarios to fully understand and optimize sensing coverage in indoor environments.
@@
This paper focuses on examining the impact of walls on sensing coverage in scenarios where all devices are located inside a room. Both theoretical analysis and experimental investigations are conducted to understand how walls influence the sensing capabilities. The results reveal that the performance of Wi-Fi sensing is significantly affected by the distance between the wall and the sensors. Fig. 1 illustrates that by reducing the distance between the wall and the Tx, the sensing coverage area can be expanded. Interestingly, the effect of walls on sensing is more intricate than previously assumed. Whether a wall enhances or reduces the sensing capability depends on the placement of the sensing devices relative to the wall. Consequently, walls can be strategically utilized to address the challenges of limited sensing coverage and enhance the robustness of the sensing system—the two most critical issues associated with wireless sensing. This study represents an important advancement towards the practical implementation of wireless sensing in real-world scenarios.
@@
  • This paper presents experimental results showcasing the practical effectiveness of our proposed model in expanding sensing coverage for stationary crowd counting applications. By strategically positioning the devices with respect to the walls, we demonstrate a notable improvement of 11.2% in the accuracy of stationary crowd counting. These empirical findings highlight the tangible benefits of applying our model, providing valuable insights for optimizing device placement and enhancing the performance of wireless sensing systems in real-world scenarios.
The rest of this paper is organized as follows. In Section 2, we provide a brief overview of the related work in the field. Section 3 presents the preliminary knowledge related to the proposed model. The model itself, along with its properties, is detailed in Section 4 and Section 5. In Section 6, we present the experimental results that validate the accuracy and usefulness of the proposed model. We discuss the limitations of our approach and potential directions for future work in Section 7 and conclude our work in Section 8.
@@ II-B Modeling sensing coverage with blocks
The aforementioned studies focus on the impact of walls on through-wall sensing scenarios, where their models necessitate that at least one of the devices or targets be located on the other side of the wall, considering only a single signal path through the wall. In contrast to these works, our research delves into wall-reflected signals with both devices and targets positioned within the indoor environment, rather than the through-wall scenario. Our objective is to investigate how the integration of wall-reflected signals and LoS signals influences the dynamic power, and hence the sensing coverage region.
@@ III-A Channel State Information
In an indoor environment, Wi-Fi signals can encounter numerous obstacles such as walls and human bodies, resulting in signal reflections and multiple paths to reach the receiver. These paths can be categorized into two types: static paths and dynamic paths, as depicted in Fig. 2. Static paths encompass the line-of-sight (LoS) path and paths reflected from stationary objects, while dynamic paths are generated by changes in target objects, such as human movements. CSI is utilized to characterize the wireless channel between a transmitter and a receiver in Wi-Fi systems [23]. The received signal can be mathematically represented by
@@ IV-B The Sensing Coverage Model in Indoor Environments Considering the Effect of Walls
where $P_{d_{wall}}$ is expressed in (12) and $P_{d_{LoS}}$ is expressed in (6). The phase difference between the two signals can be expressed as $\bigtriangleup\phi=\frac{2\pi(d_{1}+d_{2}-r_{T})}{\lambda}$, where $\lambda$ is the signal wavelength. As the static power $P_{s}$ is shown in (13), the SSNR can be expressed as
@@
Fig. 5 illustrates the heatmap of the sensing capability based on our proposed model in (18) with a Wall-Tx distance of 0.5 m. Based on our analysis, variations in the reflection coefficient do not have significant impact on the characteristics of this model; therefore, we adopt a reflection coefficient of 0.3 [32]. This model integrates the dynamic signals from LoS conditions and reflections off the wall. The heatmap employs a color scheme where yellow indicates higher sensing capability, while blue represents lower sensing capability. The rapid fluctuations between the light and dark colors are caused by the phase interactions between the two paths. However, this phenomenon is negligible compared to the dominant sensing regions near the wall and devices. Upon observing the heatmap, we notice that areas in proximity to the devices and the wall exhibit higher sensing capability. This suggests that targets located near these areas are more likely to be effectively detected.
@@ V Enhancing Sensing Coverage in Indoor Environments with Walls
In this section, we investigate the fundamental factors that impact the sensing coverage in the presence of wall reflections. We aim to understand how these factors can be effectively managed to address two prominent issues in wireless sensing: limited sensing coverage and instability due to interference. By analyzing and controlling the sensing coverage, we can mitigate the challenges posed by these issues and enhance the overall performance of wireless sensing systems.
@@ V-B Key Properties of the Indoor Sensing Coverage Model
    When the transmitter-receiver distance is small (e.g., 0.1 - 3 m in Fig. 7), the size and shape of the sensing coverage is a single small area. When the transmitter-receiver distance is large (e.g., $>$ 4 m in Fig. 7), the size and shape of the sensing coverage becomes several separated small areas surrounding the wall, transmitter, and receiver.
@@ VI Evaluation
Evaluation Setup. The key of this paper is to demonstrate the improvement contributed by the effect of walls, and the relationship between the proposed model and the performance of sensing applications in real scenarios. In the context of respiratory monitoring, our methodology integrates signal processing techniques such as the Hampel filter [34], Savitzky-Golay filter [35], and peak detection algorithms to compute respiratory rates in breaths per minute (bpm). A reference respiration rate is acquired from a Neulog Respiration Monitor Belt logger sensor NUL236. The Mean Absolute Error (MAE) metric [36] is employed to evaluate the application performance.
For stationary crowd counting accuracy evaluations, we leverage traditional machine learning models, namely Support Vector Machines (SVM), Random Forest (RF), and K-Nearest Neighbors (KNN). The preprocessing methods and parameters employed in these models remained consistent with our previous research [37]. Although more advanced signal processing and deep learning methods have the potential to achieve accuracy rates exceeding 99%, our objective is to analyze the unique enhancements provided by our sensing coverage model, which necessitates a comparative analysis against established benchmarks. Thus, we chose to utilize simple approaches to facilitate a meaningful and efficient assessment of the enhancements achieved.
@@ VI-A Evaluating Sensing Capability of the Proposed Model through Respiratory Monitoring
For the real-scenario SSNR, the dynamic power is computed based on the disparity between the CSI amplitude and the averaged amplitude within the observation window. The interference power is determined as the difference before and after filtering the CSI amplitude [17]. In Fig. 9, a respiration example is illustrated with different Wall-Tx distances. When the device is near the wall (Fig. 9a), the variance in the CSI amplitude increases, while the noise before and after filtering decreases, indicating a higher SSNR value.
@@
Experiment 1: Verify the effectiveness of using simulated SSNR to assess SSNR measured in real scenarios
In this experiment, the data with a single target respirating at various locations is examined. Fig. 10 illustrates the comparison between the simulated SSNR and the SSNR derived from real-world data. It is evident that the SSNR in the real scenario aligns with the simulated SSNR by a scale factor. This factor is contingent upon parameters such as device power, transceiver deployment, and environmental conditions. In our simulated model, we normalized the effect of these parameters, as they do not significantly impact the correlation between wall, device, target distances, and sensing coverage. These results validate the efficacy of our proposed SSNR-based model in assessing sensing capability and coverage.
Figure 10: Verify the effect of the SSNR metric with a target in various positions.
@@
In this experiment, we aim to verify the impact of the wall-device distance on the sensing capability. We position the transmitter at distances of 0.1 m, 0.5 m, and 1 m away from the wall, as depicted in Fig. 11. The distance between the Tx and Rx is fixed at 3 m, with the Rx moving simultaneously with the Tx. Fig. 12 presents the MAE (line chart) obtained for the three cases, along with the corresponding SSNR value (bar chart). We observe that the experimental result aligns well with the simulation trend. Specifically, as the distance between the wall and the device decreases within a limited range, the sensing capability increases, and the MAE decreases. This finding confirms the relationship between the wall-device distance and the extent of the sensing capability, as predicted by the simulation analysis.
Figure 12: Evaluation results of the effect of the wall-device distance without interference.
@@
The aim of this experiment is to verify the impact of the wall-device distance on sensing area expansion and its potential negative effect when interferers are present within this expanded area. The experimental setup mirrors that of Experiment 2, with the addition of an interferer within the sensing area. In Fig. 13, the green bars depict SSNR values in real scenarios, while the blue line represents the MAE for the target within the room.
As the transmitter moves closer to the wall, the SSNR decreases, indicating significant interference in this case. The trend observed in Fig. 13 demonstrates that as the wall-device distance increases, the MAE decreases. This pattern validates the adverse impact of a short wall-device distance when interferers are present within the sensing range.
Figure 13: Evaluation results of the effect of the wall-device distance with interference.

### v4

@@ I Introduction
The recent exploration of channel state information (CSI) in Wi-Fi systems has revealed the untapped potential of communication signals for diverse indoor sensing applications [1, 2]. In indoor environments, Wi-Fi signals encounter reflections from a multitude of objects, including walls and human bodies, resulting in complex multi-path propagation. Variations in CSI values predominantly arise from the dynamic movements of humans. This enables a wide range of sensing applications, including human activity recognition (HAR) [3, 4], respiration monitoring [5, 6], crowd counting [7, 8], and indoor positioning [9, 10]. While prior work has made substantial progress in enhancing CSI-based sensing performance through advanced signal processing and deep learning techniques [11, 12], comparatively little effort has focused on improving the quality of the raw measurements. In practice, the effectiveness of sensing is often limited not by post-processing but by suboptimal deployment strategies, particularly the placement of devices, which directly shapes the strength and structure of multipath signals.
One commonly overlooked aspect of deployment is the placement of devices relative to surrounding walls, an element that can significantly influence multipath propagation, but has received limited attention in existing sensing models. In practical deployments, Wi-Fi transmitters, such as routers, are typically positioned near walls to avoid obstruction and facilitate wiring. However, prior work has primarily focused on the distance between the transmitter (Tx) and receiver (Rx) as the dominant placement factor [13], while largely neglecting the impact of walls on dynamic signal paths in indoor environments. Although some studies have examined the positive role of walls in through-wall sensing scenarios [14], these models assume that at least one device is located outside the room, thereby omitting the influence of indoor wall reflections. This reveals a critical gap in current understanding: the lack of a comprehensive model that accounts for indoor wall-reflected signals under common deployment scenarios. Bridging this gap is essential for optimizing sensing coverage in real-world environments.
@@
To address this gap, we propose the first analytical model that explicitly incorporates indoor wall-reflected dynamic signal paths into Wi-Fi sensing analysis. The model focuses on the dominant components, line-of-sight (LoS) and wall-reflected paths, to capture the primary interactions that shape sensing performance. While simplified by design, it remains lightweight, generalizable, and offers tractable insights into how wall proximity influences sensing coverage. This facilitates deployment-aware system design without relying on exhaustive simulation or dense empirical datasets. We complement the theoretical model with extensive experimental investigations to understand how wall placement affects real-world sensing performance. Results show that the distance between the wall and the sensing devices significantly impacts sensing coverage. As illustrated in Fig. 1, reducing the wall-to-transmitter distance can notably expand the effective sensing area. Interestingly, while wall reflections can enhance sensing coverage within the room, they may also increase unwanted sensitivity to movements beyond the wall, depending on transceiver placement. These findings highlight the potential to strategically leverage wall proximity to address key limitations in sensing coverage and robustness. This work represents a substantial step toward practical, deployment-optimized Wi-Fi sensing in real-world indoor environments.
@@
  • This paper presents experimental results showcasing the practical effectiveness of our proposed model in expanding sensing coverage for respiratory monitoring and stationary crowd counting applications. By strategically positioning the devices with respect to the walls, we demonstrate a notable improvement of 11.2% in the accuracy of stationary crowd counting. These results confirm the practical value of our model in guiding device placement and improving sensing performance.
The rest of this paper is organized as follows. In Section II, we provide a brief overview of the related work in the field. Section III presents the preliminary knowledge related to the proposed model. The model itself, along with its properties, is detailed in Sections IV and V. In Section VI, we present the experimental results that validate the accuracy and usefulness of the proposed model. We discuss the limitations of our approach and potential directions for future work in Section VII, and conclude our work in Section VIII.
@@ II-B Modeling sensing coverage with blocks
The aforementioned studies focus on the impact of walls on through-wall sensing scenarios, where their models necessitate that at least one of the devices or targets be located on the other side of the wall, considering only a single signal path through the wall. In contrast to these works, our research delves into wall-reflected signals with both devices and targets positioned within the indoor environment, rather than the through-wall scenario. Our objective is to investigate how the integration of wall-reflected and LoS signals influences dynamic power, and thereby affects the sensing coverage region.
@@ III-A Channel State Information
In an indoor environment, Wi-Fi signals can encounter numerous obstacles such as walls and human bodies, resulting in signal reflections and multiple paths to reach the receiver. These paths can be categorized into two types: static paths and dynamic paths, as depicted in Fig. 2. Static paths encompass the LoS path and paths reflected from stationary objects, while dynamic paths are generated by changes in target objects, such as human movements. CSI is utilized to characterize the wireless channel between a transmitter and a receiver in Wi-Fi systems [19]. The received signal can be mathematically represented by
@@ IV-B The Sensing Coverage Model in Indoor Environments Considering the Effect of Walls
where $P_{d_{wall}}$ is expressed in (12) and $P_{d_{LoS}}$ is expressed in (6). The phase difference between the two signals can be expressed as $\bigtriangleup\phi=\frac{2\pi(d_{1}+d_{2}-r_{T})}{\lambda}$, where $\lambda$ is the signal wavelength.
Eq. (16) gives the combined dynamic power considering the LoS and the main-wall reflection only. However, in practical indoor environments, additional weak reflections may arise from side walls or other objects. To evaluate their influence, an error bound analysis is conducted to assess the effect of ignoring such reflections and to ensure that the simplified single-wall assumption does not alter the overall conclusions. Specifically, the additional dynamic power contributed by the side-wall reflection can be expressed as
$\Delta P$ $=P_{d_{side}}+2\sqrt{P_{d_{LoS}}P_{d_{side}}}\cos(\Delta\phi_{side})$ (17)
$+2\sqrt{P_{d_{wall}}P_{d_{side}}}\cos(\Delta\phi_{side}-\Delta\phi_{\text{wall}}),$
where $\Delta P$ denotes the side-wall contribution beyond the baseline model in (16). Here, $P_{d_{side}}$ follows the same expression as $P_{d_{wall}}$ in (12), with distances $d_{3}$ and $d_{4}$ and reflection coefficient $R_{side}$, while $\Delta\phi_{side}$ is defined as $\Delta\phi_{side}=\frac{2\pi\big((d_{3}+d_{4})-r_{T}\big)}{\lambda}$. The total dynamic power including the side-wall reflection can then be expressed as
$P_{full}=P_{d}+\Delta P.$ (18)
To quantify the approximation error introduced by ignoring the side-wall contribution, we define the relative error between the full and simplified models as
$\mathrm{RelErr}=\frac{|P_{full}-P_{d}|}{P_{full}}.$ (19)
As a concrete example, we evaluate (19) in the meeting room used for data collection (room size $6.3\,\text{m}\times 8.4\,\text{m}$), where the transmitter is placed $1\,\text{m}$ from the main wall and $3.15\,\text{m}$ from the side wall. The average relative error refers to the spatial average of $\mathrm{RelErr}$ when the human target is placed at multiple positions across the room area. To obtain an upper bound, all paths are assumed to be in phase ($\cos(\Delta\phi)=1$). Under this configuration, the average relative error is $3.26\%$ (approximately $0.14\,\text{dB}$). These results indicate that, for this practical deployment, ignoring the side-wall contribution introduces only minor deviations.
Given that the side-wall contribution is shown to be negligible, we proceed with the simplified single-wall model in the subsequent SSNR analysis. As the static power $P_{s}$ is shown in (13), the SSNR can be expressed as
@@
Fig. 5 illustrates the heatmap of the sensing capability based on our proposed model in (21), with a wall–transmitter distance of 0.5 m. The model integrates dynamic signals from both LoS conditions and wall reflections. As analyzed, variations in the reflection coefficient have limited influence on the model characteristics; thus, we adopt a fixed reflection coefficient of 0.3 [28]. Since this work primarily investigates the impact of indoor wall reflections on sensing performance, we do not explicitly model through-wall propagation. Instead, following a simplified geometric modeling perspective inspired by [14], we use a common attenuation coefficient for both reflection and (if any) penetration paths (i.e., assuming $R_{\text{wall}}=T_{\text{wall}}$). This simplification does not affect our core conclusions, as the sensing capability analysis mainly focuses on regions inside or adjacent to the wall. In the heatmap, yellow indicates higher sensing capability, while blue represents lower capability. The rapid fluctuations between light and dark regions result from phase interactions between the two paths. However, this effect is minor compared to the dominant sensing regions near the wall and transceivers. Overall, the heatmap shows that regions close to devices and the wall exhibit a higher sensing capability, suggesting that targets in these areas are more likely to be detected.
@@ V Enhancing Sensing Coverage in Indoor Environments with Walls
In this section, we investigate the fundamental factors that impact the sensing coverage in the presence of wall reflections by illustrating the theoretical sensing-coverage model derived in Section IV with numerical simulations. The analytical results presented here serve as theoretical validation before experimental verification in Section VI. Building on this theoretical analysis, we aim to understand how these factors can be effectively managed to address two prominent issues in wireless sensing: limited sensing coverage and instability due to interference. By analyzing and controlling the sensing coverage, we can mitigate the challenges posed by these issues and enhance the overall performance of wireless sensing systems.
@@ V-B Key Properties of the Indoor Sensing Coverage Model
    When the transmitter-receiver distance is small (e.g., 0.1 - 3 m in Fig. 7), the size and shape of the sensing coverage is a single small area. When the transmitter-receiver distance is large (e.g., >> 4 m in Fig. 7), the size and shape of the sensing coverage becomes several separated small areas surrounding the wall, transmitter, and receiver.
@@ VI Evaluation
Evaluation Setup. The key of this paper is to demonstrate the improvement contributed by the effect of walls, and the relationship between the proposed model and the performance of sensing applications in real scenarios. We conduct experiments in several representative indoor environments with different room layouts and wall materials to provide initial validation of the robustness and general applicability of our findings. In the context of respiratory monitoring, our methodology integrates signal processing techniques such as the Hampel filter [30], Savitzky-Golay filter [31], and peak detection algorithms to compute respiratory rates in breaths per minute (bpm). A reference respiration rate is acquired from a Neulog Respiration Monitor Belt logger sensor NUL236. The Mean Absolute Error (MAE) metric [32] is employed to evaluate the application performance, defined as
$\mathrm{MAE}=\frac{1}{M}\sum_{m=1}^{M}\left|\hat{b}_{m}-b_{\mathrm{ref}}\right|~\text{(bpm)},$ (22)
where $\hat{b}_{m}$ and $b_{\mathrm{ref}}$ denote the estimated and reference respiration rates, respectively.
For stationary crowd counting accuracy evaluations, we leverage traditional machine learning models, namely Support Vector Machines (SVM), Random Forest (RF), and K-Nearest Neighbors (KNN). The accuracy metric is defined as
$\mathrm{Acc}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}(\hat{c}_{m}=c_{m}),$ (23)
where $\mathbb{I}(\cdot)$ equals 1 when the predicted label $\hat{c}_{m}$ matches the ground truth $c_{m}$, and 0 otherwise. The preprocessing procedures and model parameters employed in these models remained consistent with our previous research [33]. Although more advanced signal processing and deep learning methods have been reported to achieve accuracy rates exceeding 99% [34], such saturation would obscure the relative differences across transceiver placements. Therefore, we intentionally adopt simple and well-validated models to ensure controlled and meaningful comparisons, allowing a clear evaluation of the sensing coverage enhancements introduced by wall proximity.
@@ VI-A Evaluating Sensing Capability of the Proposed Model through Respiratory Monitoring
For the real-scenario SSNR, the dynamic power is computed based on the disparity between the CSI amplitude and the averaged amplitude within the observation window. The interference power is determined as the difference before and after filtering the CSI amplitude [13]. As illustrated in Fig. 9, when the device is positioned close to the wall (0.1 m, Fig. 9a), the CSI amplitude exhibits a high-quality respiratory pattern with minimal noise and a peak-to-peak amplitude exceeding 100. In contrast, at a Wall–Tx distance of 1 m (Fig. 9c), the signal becomes significantly noisier and the amplitude variation drops below 20. This degradation in signal clarity and amplitude range results in a lower SSNR value, validating the advantage of wall-proximity deployment for respiration-based sensing.
@@
Experiment 1: Verify the effectiveness of using simulated SSNR to assess the SSNR measured in real scenarios
In this experiment, the data with a single target respirating at various locations is examined. Fig. 10 illustrates the comparison between the simulated SSNR and the SSNR derived from real-world data. It is evident that the SSNR in the real scenario aligns with the simulated SSNR by a scale factor. This factor is contingent upon parameters such as device power, transceiver deployment, and environmental conditions. In our simulated model, we normalized the effect of these parameters, as they do not significantly impact the correlation between wall, device, target distances, and sensing coverage. These results validate the effectiveness of our proposed SSNR-based model in assessing sensing capability and coverage.
Figure 10: Comparison between simulated and measured SSNR at different target positions. Bars: simulated SSNR; line: measured SSNR.
@@
In this experiment, we aim to verify the impact of the wall-device distance on the sensing capability. We position the transmitter at distances of 0.1 m, 0.5 m, and 1 m away from the wall, as depicted in Fig. 11. The distance between the Tx and Rx is fixed at 3 m, with the Rx moving simultaneously with the Tx. Fig. 12a presents the MAE (line chart) obtained for the three cases, along with the corresponding SSNR value (bar chart). We observe that the experimental result aligns well with the simulation trend. Specifically, as the distance between the wall and the device decreases within a limited range, the sensing capability increases, and the MAE decreases. This finding confirms the relationship between the wall-device distance and the extent of the sensing capability, as predicted by the simulation property described as Factor I in Section V-A.
Figure 12: Evaluation results of the effect of the wall–device distance on sensing performance. (a) Without interference. (b) With interference. Bars: SSNR; lines: MAE.
@@
The aim of this experiment is to verify the impact of the wall-device distance on sensing area expansion and its potential negative effect when interferers are present within this expanded area. The experimental setup mirrors that of Experiment 2, with the addition of an interferer seated approximately 0.5 meters outside the wall and capable of upper-body movement. In Fig. 12b, the green bars depict SSNR values in real scenarios, while the blue line represents the MAE for the target within the room.
As the transmitter moves closer to the wall, the SSNR decreases, indicating significant interference in this case. The trend observed in Fig. 12b demonstrates that as the wall-device distance increases, the MAE decreases. This phenomenon occurs because a shorter wall–device distance not only strengthens the main-wall reflection within the room but also increases the transmission of signals through the wall, allowing external motions to couple into the sensing channel. While the present study primarily focuses on wall-reflection effects within the room, this observation highlights that excessive proximity to the wall may also amplify cross-wall interference through transmitted components.